OpenStudy (anonymous):

Simplify the expression. [4 +(-8)] x (-7)

OpenStudy (zzr0ck3r):

\([4+(-8)]\times(-7)=[4-8]\times(-7)=(-4)\times (-7)=4 \times 7 = ?\)